jhellai20:

Write the equation of a line that has a slope 3/7 and has a y-intercept of -5. Can you guys help me find the slope int, pt-slope, and the standard form?

jhellai20:

can you show me your work and how you got these? if not, then its fine but thank you so much

jhellai20:

I have another question: Write the equation of a line that is parallel to y-3=2/3(x-9) and contains the point (-3,-4) in slope int form

Vocaloid:

First question: slope 3/7 and has a y-intercept of -5 slope-intercept form: y = mx + b where m is the slope and b is the y-intercept. plug in the appropriate m and b values. point slope: y - y1 = m(x-x1) this equation is meant to work with any potential (x1,y1) on the line, so you only need to plug in m standard form: Ax + By = C take your slope-intercept form, and re-write so that the x and y terms are on one side, and the constant is on the other. you can accomplish this by adding or subtracting terms from both sides as necessary.

Vocaloid:

Second question: Write the equation of a line that is parallel to y-3=2/3(x-9) and contains the point (-3,-4) in slope int form this is in point-slope form. the slope is 2/3, which means the equation of the line you'll be writing also has slope 2/3. (recall: parallel lines have the same slope) so far, in point slope form, we have y - y1 = (2/3)(x-x1) since they give you the point (-3,-4) you can plug this into (x1,y1) in the point-slope form. it asks for slope-int form, so solve this new equation for y.